Tom took a trip of 1,020 miles. He traveled by train at 55mph and the same number of hours by plane at 285mph. How many hours did the trip take?
I tried T=d/r 1020/55=18.545 1020/285=5.18 18.545+5.18=23.73

Found 2 solutions by rfer, stanbon:
Answer by rfer(16322)

x=time for each
55x+285x=1020
340x+1020
x=3 hrs for each
--------------------
55*3=165
285*3=855
total=1020

Answer by stanbon(75887)

Tom took a trip of 1,020 miles. He traveled by train at 55mph and the same number of hours by plane at 285mph. How many hours did the trip take?
--------------------------
Train DATA:
distance = x miles ; rate = 55 mph ; time = x/55 hrs
----------------------------------
Plane DATA:
distance = 1020-x miles ; rate = 285 mph ; time = (1020-x)/285 hrs
------------------------------------
Equation:
train time = plane time
x/55 = (1020-x)/285
Cross-multiply to get:
285x = 55(1020-x)
285x = 55*1020 - 55x
340x = 55*1020
x = 165 miles
----
train time = 165/55 = 3 hrs
plane time = (1020-165)/285 = 3 hrs
------------------------------------------
Total time = 6 hrs.
=============================================
Cheers,
Stan H.